OpenStudy (anonymous):

What is the slope of a line perpendicular to x - 5y = -10?

OpenStudy (anonymous):

? anybody care to help I got m=2 correct or missing a step?

OpenStudy (anonymous):

gradient 1 X gradient 2 = -1 for perpendicular lines. gradient of the slope you have can be found if you get the equation into y intercept form. So y to one side and x to other... x+10 = 5y so 5y/5 = x/5 + 10/5 or y = (1/5)x + 2 so m1 * m2 = -1 if m1 = 0.2, 0.2*m2 = -1 m2 = -1/.2 = -5. Test -5 * .2 = -1 A perpendicular line has slope of -5. Test y = -5x and y = 0.2x

OpenStudy (anonymous):

oh wow I did wrongggggg thanks for writing itout makes it easier to understand